OpenStudy (anonymous):

Need a little Geometry help please. I don't really understand it. The answer I get isn't shown

OpenStudy (anonymous):

ok ask away

OpenStudy (anonymous):

When I work it out \[V = \frac{ 1 }{ 3 }(10*10)(13) = 433.33\]

OpenStudy (anonymous):

what are the awnser choices?

OpenStudy (anonymous):

A. 600 ft3 B. 33.3 ft3 C. 1200 ft3 D. 400 ft3

OpenStudy (anonymous):

im doing it eright now

OpenStudy (anonymous):

Ok, Thank you.

OpenStudy (anonymous):

i got D

OpenStudy (anonymous):

you find surface area, the area of the 4 triangles, and the square's area

OpenStudy (anonymous):

let @iGreen check it though

OpenStudy (anonymous):

@iGreen cme check his work im not 100% sure im right

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

I was thinkig it was B or D so it helped

OpenStudy (igreen):

You want to find the Volume?

OpenStudy (anonymous):

the surface area

OpenStudy (anonymous):

Find the volume of the square pyramid shown. Round to the nearest tenth if necessary.

OpenStudy (igreen):

Okay

OpenStudy (igreen):

Oh, we can find the height by using the pythagorean theorem.

OpenStudy (anonymous):

ya, but this guy is in 6th grade so use small words for him

OpenStudy (anonymous):

Ok, Please don't assume I am a guy. I am a female and older then a 6th grader. I am doing home schooling to get my HS diploma and haven't been in school for years.

OpenStudy (igreen):

Lol

OpenStudy (anonymous):

oh im sorry about that, just cuz i learned this back in 6th grade

OpenStudy (igreen):



OpenStudy (anonymous):

igreen forget the mods just give him the awnser xD

OpenStudy (igreen):

I got 10 from the base of the square pyramid. Now plug in what we know into the Pythagorean Theorem. \(a^2 + b^2 = c^2\) Where 'a' an 'b' are the two legs and 'c' is the hypotenuse. \(a^2 + 5^2 = 13^2\) Simplify 5^2 and 13^2, what do you get?

OpenStudy (anonymous):

i dont remember it being that complex...back in...6th grade...

OpenStudy (igreen):

Lol, who said this was 6th grade?

OpenStudy (igreen):

@bitesize007 Could you simplify 5^2 and 13^2?

OpenStudy (anonymous):

25 and 169

OpenStudy (anonymous):

(guys behave zarkon is watching)

OpenStudy (igreen):

Yes, so we have: \(a^2 + 25 = 169\) Subtract 25 to both sides, what's 169 - 25?

OpenStudy (anonymous):

144 but that's not one of the answers

OpenStudy (anonymous):

its not done

OpenStudy (anonymous):

yet

OpenStudy (igreen):

^^^^

OpenStudy (anonymous):

well brb someone is calling me :)

OpenStudy (igreen):

Yes, so we have: \(a^2 = 144\) Find the square root of both sides: \(a = \sqrt{144}\) Find the square root of 144

OpenStudy (anonymous):

12

OpenStudy (igreen):

Yes, so the height is 12. Formula for volume of a square pyramid: \(V = \dfrac{1}{3}bh\) Plug in what we know: \(V = \dfrac{1}{3}(10)(12)\) Simplify

OpenStudy (anonymous):

43.3

OpenStudy (anonymous):

igreen remains smarter then me :3

OpenStudy (igreen):

No, I got something else, multiply again.

OpenStudy (anonymous):

i got 33.3

OpenStudy (anonymous):

40

OpenStudy (igreen):

Oh, wait that's base squared..

OpenStudy (igreen):

Multiply 1/3 * 10 * 10 * 12

OpenStudy (igreen):

Well, just multiply 10 * 40

OpenStudy (anonymous):

400

OpenStudy (anonymous):

wait so D was right :3, cool lol

OpenStudy (anonymous):

igreen made me have second thoughts xD

OpenStudy (anonymous):

Thank you very much for both of your help.

OpenStudy (igreen):

Yep, that's our answer :P

OpenStudy (igreen):

I actually just found out how to solve this right now..lol.

OpenStudy (igreen):

I guess we learn something new every day.